OpenStudy (ashleyy_smith):

Translate the sentence into an equation. "The difference of a number divided by 2 and 6 equals 8." Use the variable w for the unknown number.

OpenStudy (anonymous):

x/2 + 6 = 8 I believe

OpenStudy (anonymous):

@fay-Fay difference = subtraction

OpenStudy (ashleyy_smith):

So it's the same answer just switch the sign?

OpenStudy (anonymous):

I guess so lol that's true I forgot difference does mean subtraction

OpenStudy (ashleyy_smith):

Is that a fraction? x/2?

OpenStudy (anonymous):

No it's division

OpenStudy (ashleyy_smith):

Can you help me with one more? "Four more than the quotient of a number and 3 equals 2" Use the variable x for the unknown number/ @Fay-Fay

OpenStudy (anonymous):

Sure :) I believe it is either: 4 + x + 3 = 2 Or 4 + x/y + 3 = 2 Because the definition of quotient is "the result of division" but I honestly feel more confident in my first answer then my second, hope I helped somewhat. :)

OpenStudy (ashleyy_smith):

Got it wrong :(

Directrix (directrix):

On that first problem, the instructions specified that w should be the variable.

OpenStudy (ashleyy_smith):

it gave me a new one, "twice the difference of a number and 5 is 3" Use y for unknown number

OpenStudy (ashleyy_smith):

I think its 2-5y=3?

Directrix (directrix):

"twice the difference of a number and 5 is 3" y is the unknown number difference of y and 5 is (y -5) "twice the difference of a number and 5 is 2(y - 5) "twice the difference of a number and 5 is 3" 2 (y - 5) = 3 --> what I got

OpenStudy (anonymous):

This guy knows his stuff

Directrix (directrix):

>> 2-5y=3? not correct Look at my work above.

Directrix (directrix):

Questions? @Ashleyy_Smith

OpenStudy (anonymous):

2 X (y - 5) = 3? Try that

OpenStudy (ashleyy_smith):

ohhhhh!!

OpenStudy (anonymous):

I get it. She thinks that "and" means to add. That not how that word is being used.

OpenStudy (ashleyy_smith):

Do you know how to do this? Dale drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 7hours. When Dale drove home, there was no traffic and the trip only took 4 hours. If his average rate was 27miles per hour faster on the trip home, how far away does Dale live from the mountains?

OpenStudy (anonymous):

I am very curious now, how is the word and being used? I came in here confident but I really should have known I was going to hit my downfall lol math is not my strong point @TheDebunker

OpenStudy (anonymous):

@fay-fay "and" is being used to INCLUDE a number. The 5 more than the difference of a number and 4 equals 12 is written as 5 + (x - 5) = 12 not 5 + x/y = 12

OpenStudy (anonymous):

It's a bad example but it should give you guys an idea.

OpenStudy (anonymous):

Wait what happened to the 4 though?

Directrix (directrix):

@Fay-Fay The word "and" is being used loosely. When I read something like "What is the difference of 2 and 3," I think that the answer could be 2 - 3 = -1 or 3 - 2 = 1 just as well. These are poorly worded problems in my opinion.

OpenStudy (anonymous):

Like I said horrible example.

OpenStudy (anonymous):

When I explain it

OpenStudy (anonymous):

I see thanks guys @Directix @TheDebunker if I could give you both medals I would but this isn't even my question haha

OpenStudy (anonymous):

@Directrix poorly worded or misunderstood?

OpenStudy (anonymous):

Let's answer her last question

OpenStudy (ashleyy_smith):

(: